OpenStudy (anonymous):

The volume of a cube is 48 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

The volume of the cone is 1/3 the volume of the cube. I divide 48 by 3 and i get 16. The volume of a pyramid is 1/3 Bh, where B is the area of the base and h is the height. The volume of a cube is a^3 where a is the side of the cube. If a pyramid fits in a cube with side "a", then the base is a square, with sides a by a, and the height is also a.When you find the volume of this pyramid using the formula, you get: V = 1/3 Bh = 1/3 (a * a) * a = 1/3 a^3, which is 1/3 the volume of the cube which was a^3.

OpenStudy (anonymous):

@is3535 this is what i got but idk if i right at all im not very good at these questions

OpenStudy (anonymous):

That's correct

OpenStudy (anonymous):

@M4thM1nd thank you! can u help me with another

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 18 inches and a height of 9 inches. Pyramid B has a volume of 3136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percent. Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

Ok, we know the equation for the pyramid volume, which is: \[V = \frac{1}{3}A_{base}h\]

OpenStudy (anonymous):

What is the \(A_{base}\) of pyramid A?

OpenStudy (xapproachesinfinity):

your answer is good \[V_{pyramid}=\frac{1}{3}a^3=\frac{1}{3}V_{cube}\]

OpenStudy (xapproachesinfinity):

you reasoning is just perfect :)

OpenStudy (anonymous):

thank you! @xapproachesinfinity one sec @M4thM1nd im going to try this on my own then ppst what i got is that okay??

OpenStudy (anonymous):

Okay ^^

OpenStudy (anonymous):

umm so this is what i got so far I got the volume of Pyramid A: Pyramid A's volume is 972 because V = Base side length^2 x h/3 V = 18^2 x 9/3 = 972.

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

okay so then do i divide the smaller number by the larger number then multiply that by 100 and round if needed

OpenStudy (anonymous):

Now, we know the volume of pyramid B, 3136 in^3, what is the percentage difference?

OpenStudy (anonymous):

then i get 972/3136 = 0.30994898 x 100 = 30.99. So would it be 31% or 30%?

OpenStudy (anonymous):

that's it xD

OpenStudy (anonymous):

31%

OpenStudy (anonymous):

thank you sooo much!!! i have a few more let me finish typing out this questions answer in my quize then ill post another if you can still help me?!?!

OpenStudy (anonymous):

No problem ^^